The country of Econostan had the following expenditures last year: $600,000 in consumption


expenses, $90,000 in government purchases, and $150,000 in investment expenses. Further,


$75,000 of the products bought last year in Econostan were made in other countries and


$55,000 of the products made in Econostan were exported to other countries. What was the


gross domestic product of Econostan?



1
Expert's answer
2021-02-10T14:57:53-0500

GDP = C+I+G+ NX substituting given values, we get

GDP = 600000+ 150000+90000 + (55000 - 75000) = "\$" 820,000